OpenStudy (anonymous):

Solve the equation for x: 7x – 3y = 1.

OpenStudy (zzr0ck3r):

add the y term to both sides 7x-3y=1 7x-3y+3y = 1+3y 7x = 1+3y now devide by 7 to get the x alone (7x)/7 = (1+3y)/7 x = (1+3y)/7 = 1/7+(3y)/7